# Neural Algorithmic Reasoning for Hypergraphs with Looped Transformers

### v2

@@ C.3 Comparison
Lemma C.3 (Comparison, formal version of Lemma C.3).
@@
  • Let $C,~{}D$ be the index for the column for comparing.
@@
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(D,S_{1}),(D,S_{2})\};\\ -1&\text{if }(a,b)\in\{(C,S_{1}),(C,S_{2})\};\\ -\Omega^{-1}&\text{if }(a,b)\in\{(B_{\mathrm{global}},S_{2}),(B_{\mathrm{local}},S_{2})\};\\ 0&\text{otherwise,}\end{cases}\quad$ $(W^{(2)})_{a,b}=\begin{cases}1&\text{if }(a,b)=(E,E);\\ \Omega&\text{if }(a,b)=(S_{1},S_{1});\\ -\Omega&\text{if }(a,b)=(S_{2},S_{1});\\ 0&\text{otherwise.}\end{cases}$
$(W^{(3)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{1},S_{1})\};\\ 0&\text{otherwise,}\end{cases}\quad$ $(W^{(4)})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(S_{1},E);\\ 0&\text{otherwise.}\end{cases},$
@@ C.4 Read Scalar From Column
Lemma C.4 (Read scalar from column, formal version of Lemma C.4).
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise.}\end{cases}$
@@
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},C_{1}),(P_{2},C_{2})\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(1)}_{V})_{a,b}=\begin{cases}2&\text{if}(a,b)=(D,E)\\ 0&\text{otherwise.}\end{cases},$
@@
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)=(E,E);\\ -\Omega&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2,3)})_{a,b}=\begin{cases}1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise,}\end{cases}$
@@ C.5 Write Scalar to Column
Lemma C.5 (Write scalar to column, formal version of Lemma C.5).
@@
The key purpose is to move the information from the global variable to the local variable. The core operation of this construction is to use hardmax to select the value from a specific position. The position encoding in Definition 3.12 satisfied that $p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ for $i\neq j$. Since the MLP layer is not needed, we set all parameters to 00.
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\}\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{V})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise.}\end{cases}$
@@ C.6 Termination
Lemma C.6 (Termination, formal version of Lemma C.6).
@@
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(1)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(2)}_{K})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ 1&\text{if }(a,b)\in\{(B_{\text{local}},1),(B_{\text{local}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{Q})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ -1&\text{if }(a,b)\in\{(B_{\text{local}},1),(B_{\text{local}},2)\};\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(E,S_{1})\}\\ -1&\text{if }(a,b)=(E,S_{2});\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2,3)})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{1},S_{1}),(S_{2},S_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{2},E)\};\\ -1&\text{if }(a,b)=(S_{1},E);\\ 0&\text{otherwise.}\end{cases}$
@@ C.7 Read from Incident Matrix
Lemma C.7 (Read from incident matrix, formal version of Lemma C.7).
@@
Following from Definition 3.6, we can either employ $\psi^{(i)}(X,\widetilde{A})$ or $\psi^{(i)}(X,\widetilde{A}^{\top})$. So, reading a column and reading a row is equivalent in our setting. Considering the reading row case, we construct one attention head as follows:
$(W_{K}^{(1)},W_{Q}^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},(P_{\text{cur}})_{1}),(P_{2},(P_{\text{cur}})_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W_{V}^{(1)})_{a,b}$ =$=\begin{cases}2&\text{if }(a,b)=(B_{\text{global}},D);\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}\;(W^{(2)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(D,D);\\ 0&\text{otherwise,}\end{cases}$
Where $W^{(2)}_{K},~{}W^{(2)}_{Q}$ are constructed to make the attention matrix as identity matrix, $W^{(2)}_{V}$ is constructed to erase value. For the MLP layer, we just make it as the residential connection by setting all parameters to 00. ∎
@@ C.8 AND
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(C,S),(D,S)\};\\ -1&\text{if }(a,b)=(B_{\mathrm{global}},S);\\ 0&\text{otherwise,}\end{cases}\quad(W^{(2,3)})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S,S)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(S,E);\\ 0&\text{otherwise,}\end{cases}$
@@ C.9 Repeat AND
Lemma C.9 (Repeat AND, formal version of Lemma C.9).
@@
In this construction, we only use MLP layers. For multi-head attention layers, we just make it as the residential connection by setting all parameters to 00.
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(D,D),(C,D),(D,S_{1})\};\\ -1&\text{if }(a,b)\in\{(E,D),(B_{\text{global}},D)(B_{\text{local}},D)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(D,D),(S_{1},S_{1})\};\\ 0&\text{otherwise,}\end{cases}\quad(W^{(4)})_{a,b}=\begin{cases}1&\text{if }(a,b)=(D,D);\\ -1&\text{if }(a,b)=(S_{1},D);\\ 0&\text{otherwise.}\end{cases}$
@@ C.10 Repeat Addition
Lemma C.10 (Repeat addition, formal version of Lemma C.10).
@@
$(W^{(1)}_{K})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{Q})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)\in\left\{(B_{\text{global}},1),(B_{\text{global}},2),(B_{\text{local}},1),(B_{\text{local}},2)\right\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{V})_{a,b}$ =$=\begin{cases}1&\text{if }(a,b)=(C,D);\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}\;(W^{(2)}_{V})_{a,b}=\begin{cases}-1&\text{if }(a,b)=(C,D);\\ 0&\text{otherwise.}\end{cases}$
Where $W^{(2)}_{K},~{}W^{(2)}_{Q}$ are constructed to make the attention matrix as an identity matrix, $W^{(2)}_{V}$ is constructed to erase value. For the MLP layer, we just make it as the residential connection by setting all parameters to 00. ∎
@@
1:$\mathrm{idx}_{\mathrm{cur}},~{}\mathrm{val}_{\mathrm{cur}}\leftarrow 0,~{}0$
2:$\mathrm{idx}_{\mathrm{best}},~{}\mathrm{val}_{\mathrm{best}}\leftarrow 0,~{}\Omega$
3:$\mathrm{visit}_{\min}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
4:$\mathrm{termination}_{\min}\leftarrow 0$triangleright Initialization of minimum value
@@
11:     $\mathrm{idx}_{\mathrm{cur}}\leftarrow\mathrm{idx}_{\mathrm{cur}}+1$triangleright Increment, Lemma C.2
12:     $\mathrm{val}_{\mathrm{cur}}\leftarrow x[\mathrm{idx}_{\mathrm{cur}}]$triangleright Read scalar from column, Lemma C.4
13:     if $\mathrm{val}_{\mathrm{cur}}<\mathrm{val}_{\mathrm{best}}$ $\triangleright$ Compare, Lemma C.3
14:         $\mathrm{val}_{\mathrm{best}},~{}\mathrm{idx}_{\mathrm{best}}\leftarrow\mathrm{val}_{\mathrm{cur}},~{}\mathrm{idx}_{\mathrm{cur}}$triangleright Update variables, Lemma C.1
@@
16:     $\mathrm{visit}_{\min}[\mathrm{idx}_{\mathrm{cur}}]\leftarrow 1$triangleright Write scalar to column, Lemma C.5
17:     $\mathrm{termination}_{\min}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma C.6
@@
4:    $\mathrm{visit}[\mathrm{start}],\mathrm{termination}\leftarrow 0$triangleright Initialization of minimum value and visiting hyperedges
@@
9:                 $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\Omega$triangleright Selection, Lemma C.1
@@
15:         GetMinimumValue $(\mathrm{dists}_{\mathrm{masked}})$triangleright Get minimum value, Lemma B.1
@@
19:             $\operatorname{dist}\leftarrow\mathrm{val}_{\mathrm{best}}$triangleright Selection, Lemma C.1
@@
22:         VisitHyperdege $(A)$triangleright Degradation, Theorem 4.1
@@
25:             $\mathrm{candidates}\leftarrow\mathrm{candidates}_{\mathrm{hyperedge}}$triangleright Selection, Lemma C.1
@@
28:             $\mathrm{candidates}[i]\leftarrow\mathrm{candidates}[i]+\operatorname{dist}$triangleright Addition, Lemma C.10
@@
31:             $\mathrm{changes}[i]\leftarrow\mathrm{candidates}[i]<\mathrm{dists}[i]$triangleright Compare, Lemma C.3
@@
35:                 $\mathrm{changes}[i]\leftarrow 0$triangleright repeat AND, Lemma C.9
@@
40:                 $\mathrm{prev}[i],~{}\mathrm{dists}[i]\leftarrow\mathrm{node},\mathrm{candidates}[i]$triangleright Selection. Lemma C.1
@@
43:         $\mathrm{visit}[\mathrm{node}]\leftarrow\mathrm{visit}[\mathrm{node}]+\mathrm{termination}_{\min}$triangleright Write scalar to column, Lemma C.5
44:         $\mathrm{termination}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit})$triangleright Trigger termination, Lemma C.6
@@
1:$\mathrm{idx}_{x}\leftarrow 0$
2:$\mathrm{idx}_{y}\leftarrow 1$
3:$\mathrm{idx}_{v}\leftarrow 0$
4:$\mathrm{termination}\leftarrow 0$
5:$\mathrm{helly}\leftarrow 1$triangleright Initialization of Helly
@@
10:     $\mathrm{hyperedge}_{v}\leftarrow A[:,\mathrm{idx}_{v}]$triangleright Read column from incident matrix, Lemma C.7
@@
13:     $\mathrm{hyperedge}_{v}\leftarrow\mathrm{hyperedge}_{v}>0$triangleright Compare, Lemma C.3
14:     $\mathrm{intersection}\leftarrow\mathrm{hyperedge}_{x}\wedge\mathrm{hyperedge}_{y}\wedge\mathrm{hyperedge}_{v}$triangleright AND, Lemma C.8
@@
18:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma C.1
@@
20:     $\mathrm{idx}_{v}\leftarrow\mathrm{idx}_{v}+1$triangleright Increment, Lemma C.2
@@
22:         $\mathrm{idx}_{v}\leftarrow 1$triangleright Selection, Lemma C.1
23:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{y}+1$triangleright Selection and Increment, Lemma C.1 and Lemma C.2
@@
27:         $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{x}+1$triangleright Selection and Increment, Lemma C.1 and Lemma C.2
@@
30:         $\mathrm{termination}\leftarrow 1$triangleright Selection, Lemma C.1

### v3

@@ C.3 Comparison
Lemma C.3 (Comparison).
@@
  • Let C,DC,~D be the index for the column for comparing.
@@
$(W^{(1)})_{a,b}$ $=\begin{cases}1&\text{if }~(a,b)\in\{(E,E),(D,S_{1}),(D,S_{2})\};\\ -1&\text{if }~(a,b)\in\{(C,S_{1}),(C,S_{2})\};\\ -\Omega^{-1}&\text{if }~(a,b)\in\{(B_{\mathrm{global}},S_{2}),(B_{\mathrm{local}},S_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)})_{a,b}$ $=\begin{cases}1&\text{if }~(a,b)=(E,E);\\ \Omega&\text{if }~(a,b)=(S_{1},S_{1});\\ -\Omega&\text{if }~(a,b)=(S_{2},S_{1});\\ 0&\text{otherwise.}\end{cases}$
$(W^{(3)})_{a,b}$ $=\begin{cases}1&\text{if }~(a,b)\in\{(E,E),(S_{1},S_{1})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}-1&\text{if }~(a,b)=(E,E);\\ 1&\text{if }~(a,b)=(S_{1},E);\\ 0&\text{otherwise.}\end{cases},$
@@ C.4 Read Scalar From Column
Lemma C.4 (Read scalar from column).
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{V})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise.}\end{cases}$
@@
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},C_{1}),(P_{2},C_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{V})_{a,b}$ $=\begin{cases}2&\text{if }(a,b)=(D,E)\\ 0&\text{otherwise.}\end{cases},$
@@
$(W^{(1)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)=(E,E);\\ -\Omega&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 0&\text{otherwise,}\end{cases}$
@@ C.5 Write Scalar to Column
Lemma C.5 (Write scalar to column).
@@
The key purpose is to move the information from the global variable to the local variable. The core operation of this construction is to use hardmax to select the value from a specific position. The position encoding in Definition 3.12 satisfied that $p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ for $i\neq j$. Since the MLP layer is not needed, we set all parameters to 0.
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\}\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{V})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise.}\end{cases}$
@@ C.6 Termination
Lemma C.6 (Termination).
@@
$(W^{(1)}_{K},W^{(1)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\left\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\right\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{V})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(B_{\text{global}},E);\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(2)}_{K})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ 1&\text{if }(a,b)\in\{(B_{\text{local}},1),(B_{\text{local}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ -1&\text{if }(a,b)\in\{(B_{\text{local}},1),(B_{\text{local}},2)\};\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(1)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(E,S_{1})\}\\ -1&\text{if }(a,b)=(E,S_{2});\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{1},S_{1}),(S_{2},S_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S_{2},E)\};\\ -1&\text{if }(a,b)=(S_{1},E);\\ 0&\text{otherwise.}\end{cases}$
@@ C.7 Read from Incident Matrix
Lemma C.7 (Read from incident matrix).
@@
[…]
$(W_{K}^{(1)},W_{Q}^{(1)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},(P_{\text{cur}})_{1}),(P_{2},(P_{\text{cur}})_{2})\};\\ 0&\text{otherwise,}\end{cases}$
$(W_{V}^{(1)})_{a,b}$ $=\begin{cases}2&\text{if }(a,b)=(B_{\text{global}},D);\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{V})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(D,D);\\ 0&\text{otherwise,}\end{cases}$
Where $W^{(2)}_{K},~W^{(2)}_{Q}$ are constructed to make the attention matrix as identity matrix, $W^{(2)}_{V}$ is constructed to erase value. For the MLP layer, we just make it as the residential connection by setting all parameters to 0. ∎
@@ C.8 AND
$(W^{(1)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(C,S),(D,S)\};\\ -1&\text{if }(a,b)=(B_{\mathrm{global}},S);\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(E,E),(S,S)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(E,E);\\ 1&\text{if }(a,b)=(S,E);\\ 0&\text{otherwise,}\end{cases}$
@@ C.9 Repeat AND
Lemma C.9 (Repeat AND).
@@
In this construction, we only use MLP layers. For multi-head attention layers, we just make it as the residential connection by setting all parameters to 0.
$(W^{(1)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(D,D),(C,D),(D,S_{1})\};\\ -1&\text{if }(a,b)\in\{(E,D),(B_{\text{global}},D)(B_{\text{local}},D)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(D,D),(S_{1},S_{1})\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)=(D,D);\\ -1&\text{if }(a,b)=(S_{1},D);\\ 0&\text{otherwise.}\end{cases}$
@@ C.10 Repeat Addition
Lemma C.10 (Repeat addition).
@@
$(W^{(1)}_{K})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(B_{\text{global}},1),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\left\{(B_{\text{global}},1),(B_{\text{global}},2),(B_{\text{local}},1),(B_{\text{local}},2)\right\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(1)}_{V})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)=(C,D);\\ 0&\text{otherwise,}\end{cases}$
@@
$(W^{(2)}_{K},W^{(2)}_{Q})_{a,b}$ $=\begin{cases}1&\text{if }(a,b)\in\{(P_{1},1),(P_{2},2),(B_{\text{global}},2)\};\\ 0&\text{otherwise,}\end{cases}$
$(W^{(2)}_{V})_{a,b}$ $=\begin{cases}-1&\text{if }(a,b)=(C,D);\\ 0&\text{otherwise.}\end{cases}$
Where $W^{(2)}_{K},~W^{(2)}_{Q}$ are constructed to make the attention matrix as an identity matrix, $W^{(2)}_{V}$ is constructed to erase value. For the MLP layer, we just make it as the residential connection by setting all parameters to 0. ∎
@@
[…]
@@
11:  $\mathrm{idx}_{\mathrm{cur}}\leftarrow\mathrm{idx}_{\mathrm{cur}}+1$ $\triangleright$ Increment, Lemma C.2
12:  $\mathrm{val}_{\mathrm{cur}}\leftarrow x[\mathrm{idx}_{\mathrm{cur}}]$ $\triangleright$ Read scalar from column, Lemma C.4
[…]
14:   $\mathrm{val}_{\mathrm{best}},~\mathrm{idx}_{\mathrm{best}}\leftarrow\mathrm{val}_{\mathrm{cur}},~\mathrm{idx}_{\mathrm{cur}}$ $\triangleright$ Update variables, Lemma C.1
@@
16:  $\mathrm{visit}_{\min}[\mathrm{idx}_{\mathrm{cur}}]\leftarrow 1$ $\triangleright$ Write scalar to column, Lemma C.5
17:  $\mathrm{termination}_{\min}\leftarrow\neg(0~\mathrm{in}~\mathrm{visit}_{\min})$ $\triangleright$ Trigger termination, Lemma C.6
@@
4:  $\mathrm{visit}[\mathrm{start}],\mathrm{termination}\leftarrow 0$ $\triangleright$ Initialization of minimum value and visiting hyperedges
@@
9:      $\mathrm{dists}_{\mathrm{masked}}[i]\leftarrow\Omega$ $\triangleright$ Selection, Lemma C.1
@@
15:    GetMinimumValue $(\mathrm{dists}_{\mathrm{masked}})$ $\triangleright$ Get minimum value, Lemma B.1
@@
19:     $\operatorname{dist}\leftarrow\mathrm{val}_{\mathrm{best}}$ $\triangleright$ Selection, Lemma C.1
@@
22:    VisitHyperdege $(A)$ $\triangleright$ Degradation, Theorem 4.1
@@
25:     $\mathrm{candidates}\leftarrow\mathrm{candidates}_{\mathrm{hyperedge}}$ $\triangleright$ Selection, Lemma C.1
@@
28:     $\mathrm{candidates}[i]\leftarrow\mathrm{candidates}[i]+\operatorname{dist}$ $\triangleright$ Addition, Lemma C.10
@@
31:     $\mathrm{changes}[i]\leftarrow\mathrm{candidates}[i]<\mathrm{dists}[i]$ $\triangleright$ Compare, Lemma C.3
@@
35:      $\mathrm{changes}[i]\leftarrow 0$ $\triangleright$ repeat AND, Lemma C.9
@@
40:      $\mathrm{prev}[i],~\mathrm{dists}[i]\leftarrow\mathrm{node},\mathrm{candidates}[i]$ $\triangleright$ Selection. Lemma C.1
@@
43:    $\mathrm{visit}[\mathrm{node}]\leftarrow\mathrm{visit}[\mathrm{node}]+\mathrm{termination}_{\min}$ $\triangleright$ Write scalar to column, Lemma C.5
44:    $\mathrm{termination}\leftarrow\neg(0~\mathrm{in}~\mathrm{visit})$ $\triangleright$ Trigger termination, Lemma C.6
@@
[…]
@@
10:  $\mathrm{hyperedge}_{v}\leftarrow A[:,\mathrm{idx}_{v}]$ $\triangleright$ Read column from incident matrix, Lemma C.7
@@
13:  $\mathrm{hyperedge}_{v}\leftarrow\mathrm{hyperedge}_{v}>0$ $\triangleright$ Compare, Lemma C.3
14:  $\mathrm{intersection}\leftarrow\mathrm{hyperedge}_{x}\wedge\mathrm{hyperedge}_{y}\wedge\mathrm{hyperedge}_{v}$ $\triangleright$ AND, Lemma C.8
@@
18:   $\mathrm{termination}\leftarrow 1$ $\triangleright$ Selection, Lemma C.1
@@
20:  $\mathrm{idx}_{v}\leftarrow\mathrm{idx}_{v}+1$ $\triangleright$ Increment, Lemma C.2
@@
22:   $\mathrm{idx}_{v}\leftarrow 1$ $\triangleright$ Selection, Lemma C.1
23:   $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{y}+1$ $\triangleright$ Selection and Increment, Lemma C.1 and Lemma C.2
@@
27:   $\mathrm{idx}_{y}\leftarrow\mathrm{idx}_{x}+1$ $\triangleright$ Selection and Increment, Lemma C.1 and Lemma C.2
@@
30:   $\mathrm{termination}\leftarrow 1$ $\triangleright$ Selection, Lemma C.1